One positive number is 2/3 of the other and theor product is 24.what is the sum of the two ?
Answer by math_helper(2461)

a = (2/3)b    (1)
ab = 24       (2)

(2/3)b*b = 24
b*b = 36
b = +/-6
Discard the negative potential solution because problem statement says "positive number."
b=6 —> a=4

Ans: 4 + 6 = 10 (edited to provide the sum)

Check: (2/3)*6 = 12/3 = 4 (ok)
4*6 = 24 (ok)
——